OpenStudy (anonymous):

The rectangle shown has width x feet, and its length is five feet longer than its width. The expression which correctly shows the area of the rectangle is

OpenStudy (anonymous):

if the width =x and length =x+5 and Area=length x width then Area = (x)(x+5)